OpenStudy (anonymous):

2x1173x50= guys help please

OpenStudy (lxelle):

(2x1173) = 2346 (2346x50) = ???

OpenStudy (anonymous):

is it 8730

OpenStudy (anonymous):

lxelle

OpenStudy (anonymous):

2x1173x50 =1173x(2x50)=1173 *100=117300

OpenStudy (rational):

use commutative/associative properties of equalities to make ur life easy : axbxc = bxaxc = bx(axc)

OpenStudy (rational):

your goal is to break the wall or head

OpenStudy (anonymous):

please anybody please help me i am an junior

OpenStudy (anonymous):

2*50=?

OpenStudy (anonymous):

the wall would break before my head >.<

OpenStudy (rational):

is this applied math course @blackburnincrew

OpenStudy (anonymous):

sh!t >.< how this is applied math course ='(

OpenStudy (anonymous):

guys which grades are you guys

OpenStudy (anonymous):

9

OpenStudy (anonymous):

i am only 6 guys so try to understand

OpenStudy (anonymous):

oh :o like my nephew :( he is also have problem with this product thingy , ok answer this 2×50=؟

OpenStudy (anonymous):

100

OpenStudy (anonymous):

good , now 1173 × 100 = ؟

OpenStudy (anonymous):

is it 3419@BSwan

OpenStudy (anonymous):

:× no ok guess this 5×100 = ?

OpenStudy (anonymous):

500

OpenStudy (anonymous):

good :) and 17 x 100 =?

OpenStudy (anonymous):

1700

OpenStudy (anonymous):

343 *100

OpenStudy (anonymous):

34300

OpenStudy (anonymous):

is this* =x

OpenStudy (anonymous):

yes * means × thus any number × 100 you only add 00 to the right so 1173 ×100=؟

OpenStudy (anonymous):

=117300

OpenStudy (anonymous):

great , now back to the question 2x1173x50 you can evaluate 50×2 easily right ? so just move the 2 2x1173x50=1173x(2x50)=173x(100)=17300 got it now ?

OpenStudy (anonymous):

thanks a lot dude

OpenStudy (anonymous):

i am gone a give you the

OpenStudy (anonymous):

best response

OpenStudy (anonymous):

np :) i just wanna teach my nephew these stuff , he had problem with math so i thought if i could teach u this , the he could learned it as well :)

OpenStudy (anonymous):

its my pleasure to study with you

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:3

OpenStudy (anonymous):

ok now help with this sum which is179x4x25